(a) Briefly describe how the resistance R of a metal resistor changes with the temperature T. Sketch a functional form of R(T) and explain why.

(b) Briefly describe how the resistance R of a semiconductor resistor changes with the temperature T. Sketch a functional form of R(T) and explain why.
Answer #1

(A)

To check A is subspace of R3 or not ,let (x_1,,y_1,z_1),(x_2,y_2,z_2)\epsilon A\&\alpha \mathbb{R}

\alpha (x_1,,y_1,z_1)+(x_2,y_2,z_2)=(\alpha x_1+x_2,\alpha y_1+y_2,\alpha z_1+z_2)since

Now since (x_1,,y_1,z_1),(x_2,y_2,z_2)\epsilon A ,

x_1<y_1<z_1 \&\\ x_2<y_2<z_2___(2)

x_1<y_1<z_1 \\ \Rightarrow \\ \alpha x_1<\alpha y_1<\alpha z_1--(1)

Now add (1) and (2) ,we get

\alpha x_1+x_2<\alpha y_1+y_2<\alpha z_1+z_2

which implies that \alpha (x_1,,y_1,z_1)+(x_2,y_2,z_2)\epsilon A

hence A is subspace of R3.

(b)

To check B is subspace of R3 or not ,let (5,,y_1,z_1),(5,y_2,z_2)\epsilon A\&\alpha \mathbb{R}

now

\alpha (5,,y_1,z_1)+(5,y_2,z_2)\\ =(5\alpha +5,\alpha y_1+y_2,\alpha z_1+z_2)\\ =(5(\alpha +1),\alpha y_1+y_2,\alpha z_1+z_2)\\ =(5,\alpha y_1+y_2,\alpha z_1+z_2)only if \alpha =0

hence in genral \alpha (5,y_1,z_1)+(5,y_2,z_2)\not\epsilon A

which implies that A is not subspace of R3

Counter example:-Take

(5,1,2),(5,2,3)\epsilon A\\ then\\ (5,1,2)+(5,2,3)=(10,3,5)\not\epsilon A,hence A is not subspace of R3.

(c)

To check C is subspace of R3 or not ,let (2x,6x,9x),(2y,6y,9y)\epsilon A\&\alpha \mathbb{R}

Now \alpha (2x,6x,9x)+(2y,6y,9y)=(2(\alpha x+y),6(\alpha x+y),9(\alpha x+y))\epsilon A\Rightarrow A is subspace of R3.

(D)

To check D is subspace of R3 or not ,let (x,y,z),(x',y',z')\epsilon A\&\alpha \mathbb{R}

\alpha (x,y,z)+(x',y',z')=(\alpha x+x',\alpha y+y',\alpha z+z')

Now

-5(\alpha x+x')+4(\alpha y+y')=\alpha (-5x+4y)+(-5x'+4y')=\alpha *0+0=0\\ \&\\ -3(\alpha x+x')+2(\alpha z+z')=\alpha (-3x+2z)+(-3x'+2z')=\alpha *0+0=0\Rightarrow

\alpha (x,y,z)+(x',y',z')\epsilon D\Rightarrow D is subspace of R3.

(E)

To check E is subspace of R3 or not ,let (x,y,z),(x',y',z')\epsilon E\&\alpha \mathbb{R}

\alpha (x,y,z)+(x',y',z')=(\alpha x+x',\alpha y+y',\alpha z+z')

Now

-5(\alpha x+x')+4(\alpha y+y')+3(\alpha z+z')=(-5x+y+3z)+(-5x'+4y'+3z')=\alpha *0+0=0

which shows that E is subspace of R3.

(F)

To check F is subspace of R3 or not ,let (x,y,z),(x',y',z')\epsilon F\&\alpha \mathbb{R}

\alpha (x,y,z)+(x',y',z')=(\alpha x+x',\alpha y+y',\alpha z+z')

Now

-6(\alpha x+x')-9(\alpha y+y')+7(z+z')=\alpha (-6x-9y+7z)+(-6x'-9y'+7z')=\alpha (-2)+(-2)\neq -2\forall \alpha

which show sthat In general \alpha (x,y,z)+(x',y',z')\not \epsilon F\Rightarrow F is not subspace of R3.

Counter example

Take

(-2,0,-2)\epsilon F\because -6(-2)-9(0)+7(-2)=12-14=-2\\ (1/3,0,0)\epsilon F\because -6(1/3)-9(0)+7(0)=-2\\ but\\ (-2,0,-2)+(1/3,0,0)=(-5/3,0,-2)\not\epsilon F\because -6(-5/3)-9(0)+7(-2)=10-14=-4